OpenStudy (anonymous):

Andreas received a new book for his birthday. He read one-eighths of the book in the first week, then thirty-four more pages in the second week. If he still has one hundred sixty-two pages to read, how many total pages are in his book? one hundred forty-six one hundred seventy-two one hundred ninety-six two hundred twenty-four

OpenStudy (anonymous):

two hundred twenty-four

OpenStudy (anonymous):

224. \[(1/8)x +34+162=x\] \[196=(7/8)x\] \[x=224\]

OpenStudy (anonymous):

34 + 162 = 196 = 7/8 of number of pages in book therefore total pages = 196 * 8/7 = 224